In 7 years Maria will be twice as old as she was 3 years ago. What is her present age?
Answer by Maths68(1474)

Let
Present age of Maria = m
In 7 years means we have to add 7 to the present age of Maria
=m+7
3 years ago means we have to subtract 3 from her present age
=m-3
then
m+7=2(m-3)
m+7=2m-6
m=2m-6-7
m-2m=-13
-m=-13
Multiply by (-1) both sides
m=13
Maria is 13 years old now.